OpenStudy (anonymous):

SOLVE: √x+√x-3 =3 (roots, radicals, and complex numbers) THANKS:)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt{x}+ \sqrt{x-3}=3 }\) square both sides... \(\LARGE\color{blue}{ \bf x+2(\sqrt{x})( \sqrt{x-3})+ (x-3)=3 }\) \(\LARGE\color{blue}{ \bf x+2\sqrt{x^2-3x}+ (x-3)=3 }\) \(\LARGE\color{blue}{ \bf x+2\sqrt{x^2-3x}+ x-3=3 }\) so afr so good? (about to post more....)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf x+2\sqrt{x^2-3x}+ x=6 }\) \(\LARGE\color{blue}{ \bf 2x+2\sqrt{x^2-3x}=6 }\) \(\LARGE\color{blue}{ \bf 2(x+\sqrt{x^2-3x})=3(2) }\) \(\LARGE\color{blue}{ \bf x+\sqrt{x^2-3x}=3 }\) tell me if you are loosing me...

OpenStudy (anonymous):

no i got it

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf x+\sqrt{x^2-3x}=3 }\) \(\LARGE\color{blue}{ \bf \sqrt{x^2-3x}=3-x }\) \(\LARGE\color{blue}{ \bf x^2-3x=9-6x+x^2 }\) \(\LARGE\color{blue}{ \bf -3x=9-6x }\) \(\LARGE\color{blue}{ \bf 3x=9 }\) \(\LARGE\color{blue}{ \bf x=3 }\)

OpenStudy (solomonzelman):

My bad....

OpenStudy (solomonzelman):

I errored in first steps...

OpenStudy (anonymous):

oh really? which step?

OpenStudy (solomonzelman):